# Transversal Hamilton cycles in digraph collections

### v1

@@ 6 Concluding remarks
In this paper, we generalize Ghouila-Houri’s theorem [19] (Theorem 1.3) to transversal version, which solves a problem proposed by Chakraborti, Kim, Lee, Seo [6]. Thereby, the transversal version of Dirac’s theorem can be deduced, which was obtained by Joos and Kim [21].
An anti-directed cycle is a digraph in which the underlying graph forms a cycle and no pair of consecutive edges forms a directed path. DeBiasio and Molla [17] proved the the anti-directed Hamilton cycle is guaranteed to appear in a digraph $D$ if $\delta^{0}(D)\geq\frac{n}{2}+1$. DeBiasio, Kühn, Molla, Osthus and Taylor [16] showed that for sufficiently large $n$, every $n$-vertex digraph $D$ with $\delta^{0}(D)\geq\frac{n}{2}$ contains every orientation of a Hamilton cycle except, possibly, the anti-directed one. It would be interesting to generalize the above two results to transversal version.
An oriented graph is a digraph with no cycle of length two. Keevash, Kühn, and Osthus [23] proved the oriented version of Dirac’s theorem, which states that every $n$-vertex digraph with minimum semidegree at least $\frac{3n-4}{8}$ contains a directed Hamilton cycle. It would be interesting to consider a transversal version of this theorem.
Question 6.1.
@@
We remark that our main technique used in this paper cannot be directly generalized to the above oriented case. This is because the absorbing structure we used does not exist in the above oriented graph collections since their semi-degree is much lower than $\frac{n}{2}$. To overcome this difficulty, one has to introduce new stable conditions and build an absorption structure using Lemma 3.1. It seems that new ideas are needed to prove the tight bound in Question 6.1.
@@ Appendix A Regularity for digraph collections
Since the proof of Theorem A.4 closely follows the proof of Theorem A.3, requiring only a repeated application of the cleaning operation on $s$ hypergraphs, we omit it here for brevity.
Let $\mathcal{D}=\{D_{i}:i\in\mathcal{C}\}$ be a collection of digraphs on a common vertex set $[n]$. Based on $\mathcal{D}$, we construct the auxiliary $4$-graph $\mathscr{H}$ as follows:
$V(\mathscr{H})=[n]\cup\mathcal{C}\cup S_{1}\cup S_{2},\ \text{where}\ |S_{1}|=|S_{2}|=n,$
$E(\mathscr{H})=\left\{\{i,j,c,x\}:i,j\in[n]\ \text{with}\ ij\in E(D_{c})\ {\text{for some}}\ c\in\mathcal{C}\ \text{and}\ x\in S_{1}\ \text{if}\ i<j,\ x\in S_{2}\ \text{if}\ i>j\right\}.$
@@
By increasing $L_{0}$ and decreasing $\epsilon$, $\delta$ as necessary, we may assume that $0<\frac{1}{L_{0}}\ll\epsilon\ll\delta\ll 1$. Let $L_{1}:=\frac{24L_{0}}{\delta}$. Choose further positive constants $\epsilon^{\prime},\alpha$ satisfying $\frac{1}{L_{1}}\ll\epsilon^{\prime}\ll\alpha\ll\epsilon\ll\delta\ll 1$. Let $n_{1}$ be derived from Theorem A.3 by applying it with the parameters $(\epsilon^{\prime},L_{1},4)$. Then, $n_{2}$ is obtained from Theorem A.4 using the parameters $(\epsilon^{\prime},2,L_{0},n_{1},3)$. Let $n_{0}:=\frac{4n_{2}}{\alpha}$. By increasing $n_{0}$, we may assume that $L_{1}\ll n_{1}\ll n_{2}\ll n_{0}$. Altogether,
$0<\frac{1}{n_{0}}\ll\frac{1}{n_{2}}\ll\frac{1}{n_{1}}\ll\frac{1}{L_{1}}\ll\epsilon^{\prime}\ll\alpha\ll\epsilon\ll\delta\ll 1$
@@
Combining $(c)$, we have
${\rm deg}_{\mathscr{H}^{\prime}}(v)\geq{\rm deg}_{\mathscr{H}}(v)-(d+3\epsilon^{\prime})|U|^{3}$ (6)
for all $v\in U$. Partition each cluster $U_{i}$, $i\in[K]$, into $\frac{1}{\alpha}$ subclusters of size at most $m_{0}$ so that all but at most four subclusters of $U_{i}$ have size exactly $m_{0}$ and the property that they lie entirely within one of $V$, $\mathcal{C}$, $S_{1}$ and $S_{2}$. If a subcluster does not have this property, then add it to $U_{0}$. The new exceptional set has size at most $|U_{0}|+4\alpha m^{\prime}K$, and let $V_{0}$ be its intersection with $V$, $\mathcal{C}_{0}$ be its intersection with $\mathcal{C}$. Relabel the subclusters so that those which are subsets of $V$ are $V_{1},\ldots,V_{L}$ and those which are subsets of $\mathcal{C}$ are $\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$. For each $c\in\mathcal{C}$, let $D_{c}^{\prime}$ be the spanning sub-digraph of $D_{c}$ with vertex set $V=:[n]$ and edge set
@@
We claim that the properties of the Lemma 2.2 (i)-(iv) are satisfied.
@@
$L+M\leq\frac{2K}{\alpha}\leq\frac{2n_{1}}{\alpha}\leq\frac{n_{0}}{2},$ (8)
@@
for all $v\in V$, and similarly for $\sum_{c\in\mathcal{C}}d^{-}_{D_{c}^{\prime}}(v)$. Additionally, observe that $e(D_{c})=\frac{{\rm deg}_{\mathscr{H}}(c)}{n}$ for all $c\in\mathcal{C}$. It follows that
$e(D_{c}^{\prime})\geq\frac{{\rm deg}_{\mathscr{H}^{\prime}}(c)}{n}\geq\frac{{\rm deg}_{\mathscr{H}}(c)}{n}-\left(\frac{8d}{\delta^{3}}+\frac{\epsilon}{4}\right)n^{2}=e(D_{c})-\left(\frac{8d}{\delta^{3}}+\frac{\epsilon}{4}\right)n^{2}$ (12)
for all $c\in\mathcal{C}$.
@@
Notice that we can view partition $P=(V_{1},\ldots,V_{L},\mathcal{C}_{1},\ldots,\mathcal{C}_{M})$ as comprising two different partitions $P_{1}$, $P_{2}$ of hypergraphs $\mathscr{H}_{1}$, $\mathscr{H}_{2}$ (both partitions divide $V$ and $\mathcal{C}$ in the same way, but the sets of edges among the partition sets are different). By increasing $n_{2}$, we may assume that $n_{2}\gg N(\epsilon^{\prime},2,L_{0},L+M,3)$ (which is defined in Theorem A.4). Consequently, $n_{0}\gg N(\epsilon^{\prime},2,L_{0},L+M,3)$ as well. Applying Theorem A.4, we can find a refinement partition of $P$, say $U_{0}^{\prime},U_{1}^{\prime},\ldots,U_{K^{\prime}}^{\prime}$, and spanning subhypergraphs $\mathscr{H}_{1}^{\prime}\subseteq\mathscr{H}_{1}$ and $\mathscr{H}_{2}^{\prime}\subseteq\mathscr{H}_{2}$ such that the following hold:
@@ Appendix B Transversal perfect matching in bipartite graph collections
We first add vertices with small degree to $C_{1}\cup C_{2}$. Let $\mu=\epsilon^{5}$. Since $G$ is $\epsilon$-extremal, there are $X\subseteq V_{1},Y\subseteq V_{2}$ each of size at least $\left(\frac{1}{2}-\mu\right)n$ such that $e_{G}(X,Y)\leq\mu n^{2}$. Thus, it is easy to see that for all but at most $\sqrt{\mu}n$ vertices $v$ in $X$ we have $d_{G}(v,Y)\leq\sqrt{\mu}n$. Similarly, for all but at most $\sqrt{\mu}n$ vertices $v$ in $Y$ we have $d_{G}(v,X)\leq\sqrt{\mu}n$. We add these exceptional vertices from $X$ and from $Y$ to $C_{1}$ and $C_{2}$, respectively. Now, the new sets $X$ and $Y$ each has size at least $\left(\frac{1}{2}-\mu\right)n-\sqrt{\mu}n\geq\left(\frac{1}{2}-2\sqrt{\mu}\right)n.$ Choose $X_{1}\subseteq X,Y_{1}\subseteq Y$ such that $|X_{1}|=|Y_{1}|=\left(\frac{1}{2}-2\sqrt{\mu}\right)n$.
@@
Since the number of edges between $X_{1}$ and $Z_{2}$ is at least $(\frac{1}{2}-2\sqrt{\mu})n\times\left(\frac{1}{2}-8\sqrt{\mu}\right)n$ and $|Z_{2}|=(\frac{1}{2}-2\sqrt{\mu})n$, there are at most $50\mu^{\frac{1}{4}}n$ vertices $v$ in $Z_{2}$ such that $d_{G}(v,X_{1})\leq(\frac{1}{2}-10\mu^{\frac{1}{4}})n$. Similarly, for all but at most $50\mu^{\frac{1}{4}}n$ vertices $v$ in $Z_{1}$ we have $d_{G}(v,Y_{1})\geq(\frac{1}{2}-10\mu^{\frac{1}{4}})n$. Thus, we can choose $A_{1}\subseteq X_{1}$, $B_{1}\subseteq Z_{1}$, $A_{2}\subseteq Y_{1}$, $B_{2}\subseteq Z_{2}$ such that $|A_{1}|=|B_{1}|=|A_{2}|=|B_{2}|=(\frac{1}{2}-\epsilon)n$ and $d_{G}(v,X_{i})\geq(\frac{1}{2}-2\epsilon)n$ for $v\in X_{3-i}$ and $X\in\{A,B\}$, $i\in[2]$. Let $C_{1}=V_{1}\setminus(A_{1}\cup B_{1})$ and $C_{2}=V_{2}\setminus(A_{2}\cup B_{2})$. Hence, $(A_{1},B_{1},C_{1},A_{2},B_{2},C_{2})$ is the desired partition. ∎
@@
In this subsection, we prove Theorem 2.10. Before proceeding with the proof, we first present some key lemmas that will be used in the proof.
@@
Suppose that $0<\frac{1}{n}\ll\mu^{\prime}\ll\alpha\ll\gamma,\epsilon,\ll\delta\ll 1$. Let $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ be a collection of bipartite graphs on a common vertex partition $V=V_{1}\cup V_{2}$ with $|V_{1}|=|V_{2}|=n$. If $\mathcal{G}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable, then $\mathcal{G}$ is $\mu^{\prime}$-nice.
@@
For the sake of contradiction, suppose that $\mathcal{G}$ is not $\mu^{\prime}$-nice. Then there is $A\subseteq V_{1}$ of size $\lfloor\frac{n}{2}\rfloor$ and $B\subseteq V_{2}$ of size $\lfloor\frac{n}{2}\rfloor$ such that $e_{\mathcal{G}}(A,B)\leq\mu^{\prime}n^{3}$. Therefore, for all but at most $\sqrt{\mu^{\prime}}n$ colors $i\in[n]$, we have $e_{G_{i}}(A,B)<\sqrt{\mu^{\prime}}n^{2}$, implying that such $G_{i}$ is not $\sqrt{\mu^{\prime}}$-nice. This further implies that $\mathcal{G}$ is not $(1-\sqrt{\mu^{\prime}},\sqrt{\mu^{\prime}})$-strong stable and hence not $(\gamma,\alpha)$-strong stable, since $\mu^{\prime}\ll\gamma,\alpha$.
@@
$|A\Delta A_{1}^{i}|\leq\frac{\delta n}{2}\,\,\,\text{and}\,\,\,|B\Delta B_{2}^{i}|\leq\frac{\delta n}{2}.$
Otherwise, $e_{\mathcal{G}}(A,B)\geq\frac{\delta^{3}n^{3}}{8}>\mu^{\prime}n^{3}$ by $\mu^{\prime}\ll\epsilon\ll\delta$, a contradiction. Delete such colors from $[(1-\sqrt{\mu^{\prime}})n]$ and let $\mathcal{C}$ be the remaining color set. Thus, for every $i,j\in\mathcal{C}$ the graphs $G_{i}$ and $G_{j}$ are not $\delta$-crossing. Therefore,
@@
Moreover, for each $c\in\mathcal{C}$, we have $e_{G_{c}}(N_{1}^{+},N_{2}^{+})=0$ by the maximality of $\mathcal{M}$. Thus, it suffices to show that for all $c\in{\rm col}(\mathcal{M}[N_{1},N_{1}^{+}]\cup\mathcal{M}[N_{2},N_{2}^{+}])$, we have $e_{G_{c}}(N_{1}^{+},N_{2}^{+})\leq 4\mu n^{2}$. If so,
@@
a contradiction to $\mathcal{G}$ being $6\mu$-nice. Next, we consider the case $c\in{\rm col}(\mathcal{M}[N_{1},N_{1}^{+}])$; the proof for the case $c\in{\rm col}(\mathcal{M}[N_{2},N_{2}^{+}])$ is similar and is omitted.
@@
$|N_{3}^{+}\Delta N_{1}^{+}|=|N_{3}^{+}\cup N_{1}^{+}|-|N_{3}^{+}\cap N_{1}^{+}|\leq(n-|N_{2}|)-(|N_{1}|+|N_{3}|-(n-|N_{2}|))\leq 4\mu n.$ (13)
On the other hand, for the maximality of matching $\mathcal{M}^{\prime}$, we have $e_{G_{c}}(N_{3}^{+},N_{2}^{+})=0$. Thus, together with (13), we have $e_{G_{c}}(N_{1}^{+},N_{2}^{+})\leq 4\mu n^{2}$, as desired. ∎
@@
We define a vertex $v\in V$ is $i$-good if either $G_{i}$ is $\epsilon$-extremal and $v\in V\setminus(C^{1}_{i}\cup C^{2}_{i})$ or $G_{i}$ is not $\epsilon$-extremal and $d_{G_{i}}(v)\geq(\frac{1}{2}-\epsilon^{3})n$. Note that for each $i\in[n]$ and $j\in\{1,2\}$, there are at most $2\epsilon n$ vertices in $V_{j}$ that are not $i$-good.
@@
Without loss of generality, we assume that $G_{1},\ldots,G_{\gamma n}$ are $\alpha$-nice. For $i\in[\lambda n]$, define
@@
$Z_{i}(c,uv):=\{(w,w^{\prime})\in\mathcal{F}_{i}:ww^{\prime}\,\,\text{is a}\,\,c\text{-absorbing edge of}\,\,(u,v)\}\ \text{and}\ Z(c,uv):=\bigcup_{i\in[\lambda n]}Z_{i}(c,uv).$
@@
[…]
@@
Let $10<\frac{1}{n}\ll d\ll\lambda,\mu\ll\epsilon\ll\delta<$ 1. Suppose that $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ is a collection of bipartite graphs on a common vertex partition $V=V_{1}\cup V_{2}$ with $|V_{1}|=|V_{2}|=n$ such that the following conditions hold:
@@
Fix such a $\mathcal{U}$ and let $\overline{\mathcal{U}}:=V(H)\setminus\mathcal{U}$. By relabeling colors, one may assume $\mathcal{U}=:[t]$. Clearly,
@@
[…]
@@
$\bullet$ For vertices in $V_{2}$, there are at most $2\epsilon n+dn$ vertices $w$ that are either not $c_{i}^{2}$-good or $d_{G_{c}}(w)<\left(\frac{1}{2}-\mu\right)n$. Thus, we have at least $n-2\epsilon n-dn-2\geq\frac{n}{2}$ choices for $w_{i}^{2^{\prime}}$ such that $w_{i}^{2^{\prime}}$ is $c_{i}^{2}$-good and $d_{G_{c}}(w)\geq\left(\frac{1}{2}-\mu\right)n$.
@@
$\mathcal{F}_{i}:=\{(w_{i}^{1},w_{i}^{1^{\prime}},w_{i}^{2},w_{i}^{2^{\prime}})\in V_{1}\times V_{2}\times V_{1}\times V_{2}:(c_{i}^{1},w_{i}^{1},w_{i}^{1^{\prime}})\,\,\text{and}\,\,(c_{i}^{2},w_{i}^{2},w_{i}^{2^{\prime}})\,\,\,\text{are absorbable}\}.$

### v2

@@ 8 Concluding remarks
In this paper, we establish a transversal analogue of Ghouila-Houri’s theorem [19] (Theorem 1.3), thereby resolving a problem proposed by Chakraborti, Kim, Lee, and Seo [6] for all sufficiently large $n$. As a consequence, we recover the transversal version of Dirac’s theorem, previously obtained by Joos and Kim [21].
An anti-directed cycle is a digraph in which the underlying graph forms a cycle, and no pair of consecutive edges forms a directed path. DeBiasio and Molla [17] proved that the anti-directed Hamilton cycle is guaranteed to appear in a digraph $D$ if $\delta^{0}(D)\geq\frac{n}{2}+1$. DeBiasio, Kühn, Molla, Osthus, and Taylor [16] showed that for sufficiently large $n$, every $n$-vertex digraph $D$ with $\delta^{0}(D)\geq\frac{n}{2}$ contains every orientation of a Hamilton cycle except, possibly, the anti-directed one. It would be interesting to generalize the above two results to a transversal version.
An oriented graph is a digraph with no cycle of length two. Keevash, Kühn, and Osthus [23] proved the oriented version of Dirac’s theorem, which states that every $n$-vertex digraph with minimum semi-degree at least $\frac{3n-4}{8}$ contains a directed Hamilton cycle. It would be interesting to consider a transversal version of this theorem.
Question 8.1.
@@
We remark that our main technique used in this paper cannot be directly generalized to the above oriented case. This is because the absorbing structure we used does not exist in the above oriented graph collections since their semi-degree is much lower than $\frac{n}{2}$. To overcome this difficulty, one has to introduce new stable conditions and build an absorption structure using Lemma 3.4. It seems that new ideas are needed to prove the tight bound in Question 8.1.
9 Acknowledgment
We would like to express our gratitude to the anonymous reviewers for their valuable comments that greatly improved the presentation of this paper.
@@ Appendix A Regularity for digraph collections
Since the proof of Theorem A.4 closely follows that of Theorem A.3, requiring only a repeated application of the cleaning operation on $s$ hypergraphs, we omit it here for brevity.
Let $\mathcal{D}=\{D_{i}:i\in\mathcal{C}\}$ be a collection of digraphs on a common vertex set $[n]$. Based on $\mathcal{D}$, we construct the auxiliary $4$-graph $\mathscr{H}$ as follows. The vertex set is $V(\mathscr{H})=[n]\cup\mathcal{C}\cup S_{1}\cup S_{2}$, where $S_{1}$ and $S_{2}$ are new sets with $|S_{1}|=|S_{2}|=n$, and the four sets $[n]$, $\mathcal{C}$, $S_{1}$, $S_{2}$ are pairwise disjoint. The edge set of $\mathscr{H}$ is defined by
${E(\mathscr{H})=\left\{\{i,j,c,x\}:i,j\in[n],\ c\in\mathcal{C},\ ij\in E(D_{c});\ x\in S_{1}\ \text{if}\ i<j\ \text{and}\ x\in S_{2}\ \text{otherwise}\right\}}.$
@@
By increasing $L_{0}$ and decreasing $\epsilon$, $\delta$ as necessary, we may assume that $0<\frac{1}{L_{0}}\ll\epsilon\ll\delta\ll 1$. Let $L_{1}:=\frac{24L_{0}}{\delta}$. Choose further positive constants $\epsilon^{\prime},\alpha$ satisfying $\frac{1}{L_{1}}\ll\epsilon^{\prime}\ll\alpha\ll\epsilon\ll\delta\ll 1$. Let $n_{1}$ be derived from Theorem A.3 by applying it with the parameters $(\epsilon^{\prime},L_{1},4)$. Next, let $n_{2}$ be obtained from Theorem A.4 using the parameters $(\epsilon^{\prime},2,L_{0},n_{1},3)$. Define $n_{0}:=\frac{4n_{2}}{\alpha}$. By increasing $n_{0}$ if necessary, we may assume that $L_{1}\ll n_{1}\ll n_{2}\ll n_{0}$. Altogether,
$0<\frac{1}{n_{0}}\ll\frac{1}{n_{2}}\ll\frac{1}{n_{1}}\ll\frac{1}{L_{1}}\ll\epsilon^{\prime}\ll\alpha\ll\epsilon\ll\delta\ll 1.$
@@
Combining $(c)$, for each $v\in U$ we have
${\rm deg}_{\mathscr{H}^{\prime}}(v)\geq{\rm deg}_{\mathscr{H}}(v)-(d+3\epsilon^{\prime})|U|^{3}.$
Partition each cluster $U_{i}$, $i\in[K]$, into $\frac{1}{\alpha}$ subclusters of size at most $m_{0}$ so that all but at most $4$ subclusters of $U_{i}$ have size exactly $m_{0}$ and the property that they lie entirely within one of $V$, $\mathcal{C}$, $S_{1}$ and $S_{2}$. If a subcluster does not have this property, then add it to $U_{0}$. The new exceptional set has size at most $|U_{0}|+4\alpha m^{\prime}K$, and let $V_{0}$ be its intersection with $V$, $\mathcal{C}_{0}$ be its intersection with $\mathcal{C}$. Relabel the subclusters so that those which are subsets of $V$ are $V_{1},\ldots,V_{L}$ and those which are subsets of $\mathcal{C}$ are $\mathcal{C}_{1},\ldots,\mathcal{C}_{M}$. For each $c\in\mathcal{C}$, let $D_{c}^{\prime}$ be the spanning sub-digraph of $D_{c}$ with vertex set $V=:[n]$ and edge set
@@
We claim that the properties of Lemma 2.2 (i)-(iv) are satisfied.
@@
$L+M\leq\frac{2K}{\alpha}\leq\frac{2n_{1}}{\alpha}\leq\frac{n_{0}}{2},$
@@
for all $v\in V$, and similarly for $\sum_{c\in\mathcal{C}}d^{-}_{D_{c}^{\prime}}(v)$. Additionally, observe that $e(D_{c})=\frac{{\rm deg}_{\mathscr{H}}(c)}{n}$ for each $c\in\mathcal{C}$. It follows that for each $c\in\mathcal{C}$,
$e(D_{c}^{\prime})\geq\frac{{\rm deg}_{\mathscr{H}^{\prime}}(c)}{n}\geq\frac{{\rm deg}_{\mathscr{H}}(c)}{n}-\left(\frac{8d}{\delta^{3}}+\frac{\epsilon}{4}\right)n^{2}=e(D_{c})-\left(\frac{8d}{\delta^{3}}+\frac{\epsilon}{4}\right)n^{2}.$ (9)
@@
[…]
@@ Appendix B Transversal perfect matching in bipartite graph collections
In this subsection we prove Lemma 2.7, which provides a characteristic partition of each balanced bipartite graph whose minimum degree is just below $\frac{n}{2}$, allowing a tiny fraction of vertices that may have even smaller degrees.
@@
We first add vertices with degree at most $\left(\frac{1}{2}-\epsilon^{5}\right)n$ to $C_{1}\cup C_{2}$. Let $\mu=\epsilon^{5}$. Since $G$ is $\epsilon$-extremal, there are $X\subseteq V_{1},\,Y\subseteq V_{2}$ each of size at least $\left(\frac{1}{2}-\mu\right)n$ such that $e_{G}(X,Y)\leq\mu n^{2}$. Thus, it is easy to see that for all but at most $\sqrt{\mu}n$ vertices $v$ in $X$ we have $d_{G}(v,Y)\leq\sqrt{\mu}n$. Similarly, for all but at most $\sqrt{\mu}n$ vertices $v$ in $Y$ we have $d_{G}(v,X)\leq\sqrt{\mu}n$. We add these exceptional vertices from $X$ and from $Y$ to $C_{1}$ and $C_{2}$, respectively. Now, the new sets $X$ and $Y$ each has size at least $\left(\frac{1}{2}-\mu\right)n-\sqrt{\mu}n\geq\left(\frac{1}{2}-2\sqrt{\mu}\right)n.$ Choose $X_{1}\subseteq X,Y_{1}\subseteq Y$ such that $|X_{1}|=|Y_{1}|=\left(\frac{1}{2}-2\sqrt{\mu}\right)n$.
@@
Since the number of edges between $X_{1}$ and $Z_{2}$ is at least $(\frac{1}{2}-2\sqrt{\mu})n\times\left(\frac{1}{2}-8\sqrt{\mu}\right)n$ and $|Z_{2}|=(\frac{1}{2}-2\sqrt{\mu})n$, there are at most $50\mu^{\frac{1}{4}}n$ vertices $v$ in $Z_{2}$ such that $d_{G}(v,X_{1})\leq(\frac{1}{2}-10\mu^{\frac{1}{4}})n$. Similarly, for all but at most $50\mu^{\frac{1}{4}}n$ vertices $v$ in $Z_{1}$ we have $d_{G}(v,Y_{1})\geq(\frac{1}{2}-10\mu^{\frac{1}{4}})n$. Thus, we can choose $A_{1}\subseteq X_{1}$, $B_{1}\subseteq Z_{1}$, $A_{2}\subseteq Y_{1}$, $B_{2}\subseteq Z_{2}$ such that $|A_{1}|=|B_{1}|=|A_{2}|=|B_{2}|=(\frac{1}{2}-\epsilon)n$ and $d_{G}(v,X_{i})\geq(\frac{1}{2}-2\epsilon)n$ for $v\in X_{3-i}$ and $X\in\{A,B\}$, $i\in[2]$. Let $C_{1}=V_{1}\setminus(A_{1}\cup B_{1})$ and $C_{2}=V_{2}\setminus(A_{2}\cup B_{2})$. Hence, $(A_{1},B_{1},C_{1},A_{2},B_{2},C_{2})$ is the desired partition of $G$. ∎
@@
In this subsection we prove Theorem 2.10, giving a sufficient condition for the existence of a transversal perfect matching in bipartite graph collections. Before proceeding with the proof, we first present some key lemmas that will be used in the proof.
@@
Suppose that $0<\frac{1}{n}\ll\mu^{\prime}\ll\alpha\ll\gamma,{\epsilon}\ll\delta\ll 1$. Let $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ be a collection of bipartite graphs on a common vertex partition $V=V_{1}\cup V_{2}$ with $|V_{1}|=|V_{2}|=n$. If $\mathcal{G}$ is $(\gamma,\alpha,\epsilon,\delta)$-stable, then $\mathcal{G}$ is $\mu^{\prime}$-nice.
@@
For the sake of contradiction, suppose that $\mathcal{G}$ is not $\mu^{\prime}$-nice. Then there are sets $A\subseteq V_{1}$ and $B\subseteq V_{2}$, each of size $\lfloor\frac{n}{2}\rfloor$, such that $e_{\mathcal{G}}(A,B)\leq\mu^{\prime}n^{3}$. Therefore, for all but at most $\sqrt{\mu^{\prime}}n$ colors $i\in[n]$, we have $e_{G_{i}}(A,B)<\sqrt{\mu^{\prime}}n^{2}$, implying that such $G_{i}$ is not $\sqrt{\mu^{\prime}}$-nice. This further implies that $\mathcal{G}$ is not $(1-\sqrt{\mu^{\prime}},\sqrt{\mu^{\prime}})$-strongly stable and hence not $(\gamma,\alpha)$-strongly stable, since $\mu^{\prime}\ll\gamma,\alpha$.
@@
$|A\triangle A_{1}^{i}|\leq\frac{\delta n}{2}\ \ \text{and}\ \ |B\triangle B_{2}^{i}|\leq\frac{\delta n}{2}.$
Otherwise, $e_{\mathcal{G}}(A,B)\geq\frac{\delta^{3}n^{3}}{8}>\mu^{\prime}n^{3}$ by $\mu^{\prime}\ll\epsilon\ll\delta$, a contradiction. Delete such colors from $(1-\sqrt{\mu^{\prime}})n]$ and let $\mathcal{C}$ be the remaining color set. Thus, for every $i,j\in\mathcal{C}$ the graphs $G_{i}$ and $G_{j}$ are not $\delta$-crossing. Therefore,
@@
Moreover, for each $c\in\mathcal{C}$, we have $e_{G_{c}}(N_{1}^{+},N_{2}^{+})=0$ by the maximality of $\mathcal{M}$. Thus, it suffices to show that for all $c\in{\rm col}(\mathcal{M}[N_{1},N_{1}^{+}]\cup\mathcal{M}[N_{2},N_{2}^{+}])$, we have $e_{G_{c}}(N_{1}^{+},N_{2}^{+})\leq 4\mu n^{2}$. If this holds, then
@@
contradicting the assumption that $\mathcal{G}$ is $6\mu$-nice. Next, we consider the case $c\in{\rm col}(\mathcal{M}[N_{1},N_{1}^{+}])$; the proof for the case $c\in{\rm col}(\mathcal{M}[N_{2},N_{2}^{+}])$ is similar and is omitted.
@@
$|N_{3}^{+}\triangle N_{1}^{+}|=|N_{3}^{+}\cup N_{1}^{+}|-|N_{3}^{+}\cap N_{1}^{+}|\leq(n-|N_{2}|)-(|N_{1}|+|N_{3}|-(n-|N_{2}|))\leq 4\mu n.$ (10)
On the other hand, by the maximality of $\mathcal{M}^{\prime}$, we have $e_{G_{c}}(N_{3}^{+},N_{2}^{+})=0$. Thus, together with (10), we have $e_{G_{c}}(N_{1}^{+},N_{2}^{+})\leq 4\mu n^{2}$, as desired. ∎
@@
A vertex $v\in V$ is called $i$-good if either $G_{i}$ is $\epsilon$-extremal and $v\in V\setminus(C^{1}_{i}\cup C^{2}_{i})$ or $G_{i}$ is not $\epsilon$-extremal and $d_{G_{i}}(v)\geq(\frac{1}{2}-\epsilon^{3})n$. Note that for each $i\in[n]$ and $j\in\{1,2\}$, there are at most $2\epsilon n$ vertices in $V_{j}$ that are not $i$-good.
@@
Without loss of generality, assume that $G_{1},\ldots,G_{\gamma n}$ are $\alpha$-nice. For each $i\in[\lambda n]$, define
@@
$Z_{i}(c,uv):=\big\{(w,w^{\prime})\in\mathcal{F}_{i}:ww^{\prime}\,\,\text{is a}\,\,c\text{-absorbing edge of}\,\,(u,v)\big\}\ \ \text{and}\ \ Z(c,uv):=\bigcup_{i\in[\lambda n]}Z_{i}(c,uv).$
@@
$\mathbf{H}:=\{\mathcal{F}_{i}:i\in[\lambda n]\},\,\,\,\mathbf{Z}:=\{Z(c,uv):c\in[n],(u,v)\in V_{1}\times V_{2}\,\,\text{and}\,\,v\,\,\text{is}\,\,c\textrm{-good}\},$
@@
Let $0<\frac{1}{n}\ll d\ll\lambda,\mu\ll\epsilon\ll\delta<1$. Suppose that $\mathcal{G}=\{G_{1},\ldots,G_{n}\}$ is a collection of bipartite graphs on a common vertex partition $V=V_{1}\cup V_{2}$ with $|V_{1}|=|V_{2}|=n$ such that the following conditions hold:
@@
Fix such a set $\mathcal{U}$ and let $\overline{\mathcal{U}}:=V(H)\setminus\mathcal{U}$. By relabeling colors, one may assume $\mathcal{U}=:[t]$. Clearly,
@@
$\mathbf{Z}:=\left\{Z(i,uv):=\bigcup_{j\in\mathcal{C}_{u}\cap N_{H}(i,\mathcal{U})}Z_{j}(i,uv):i\in\overline{\mathcal{U}},u\in\overline{T_{i}}\ \text{and}\ v\ \text{is}\ i\text{-good}\right\}.$
@@
$\bullet$ For vertices in $V_{2}$, there are at most $2\epsilon n+dn$ vertices $w$ that are either not $c_{i}^{2}$-good or $d_{G_{c}}(w)<\left(\frac{1}{2}-\mu\right)n$. Thus, we have at least $n-2\epsilon n-dn-2\geq\frac{n}{2}$ choices for $w_{i}^{2^{\prime}}$ such that $w_{i}^{2^{\prime}}$ is $c_{i}^{2}$-good and $d_{G_{c}}(w_{i}^{2^{\prime}})\geq\left(\frac{1}{2}-\mu\right)n$.
@@
$\mathcal{F}_{i}:=\{(w_{i}^{1},w_{i}^{1^{\prime}},w_{i}^{2},w_{i}^{2^{\prime}})\in V_{1}\times V_{2}\times V_{1}\times V_{2}:(c_{i}^{1},w_{i}^{1},w_{i}^{1^{\prime}})\penalty 10000\ \text{and}\penalty 10000\ (c_{i}^{2},w_{i}^{2},w_{i}^{2^{\prime}})\penalty 10000\ \text{are absorbable}\}.$